OpenStudy (anonymous):

1. If you weigh 490N on earth how much do you weigh on the moon? 2. What is the gravitational force between two masses of 15kg each, when their centers are 0.25m? Could you detect this force with even sensitive equipment? 3. What would be the weight of the moon if it were resting on the surface of the earth? Remember that r is the distance between the centers of the objects. rearth = 6.37x106m mearth =5.98x1024kg. 4. What is the mass of an object that weighs 55N on earth?

OpenStudy (anonymous):

1. The acceleration in freefall on the Moon is about 1.622 m/s^2. The acceleration in freefall on the Earth is 9.8 m/s^2. First find your weight on earth: $$F = ma = mg$$ $$490 N = m(9.8 m/s^2)$$ $$\frac{490 N}{(9.8 m/s^2)} = m$$ $$50kg = m$$ Now find your weight on the moon: $$F = mg$$ $$F = (50 kg)(1.622 m/s^2)$$ $$F = 81.1N$$ Intuitively this answer makes sense, because you should weight significantly less on the moon.